OpenStudy (anonymous):

Find the part using multiplication shortcut. 55% of 1540 military personnel.

OpenStudy (anonymous):

1540*.55

OpenStudy (anonymous):

You could do it using this shortcut in your head if you notice 1540 * .55 = 1540 * 0.5 + 1540 * 0.05 = Easy to halve 1540 in your head..and for 1540 * .05 just move the decimal one place 770 + 77 = 847